# Antiferromagnetic pseudogap in the two-dimensional Hubbard model deep in the renormalized classical regime

### v1

@@ III.1 Comparison with DiagMC results
The TPSC+ study in the Ref. [gauvin2023improved] was limited to temperatures $T\geq 0.79$. In the current study, we were able to resolve convergence problem at low temperatures using methods that are explained in Appendix A. We focus, at half-filling, on the antiferromagnetic pseudogap regime and on estimates of the zero-temperature gap. We choose $U=2$ because this allows us to make comparison to the DiagMC results.
III.1.1 Double Occupancy
@@
Figure 1 shows double occupancy $D$ as function of temperature $T$. The results are for TPSC+, TPSC+Classical and for the DiagMC [Schaefer2021] benchmark results. The factor $\tilde{g}_{fb}$ for the feedback self-energy used in the calculation of the partially-dressed susceptibility $\chi^{(2)}$ was set to $\tilde{g}_{fb}=3/8$. Both TPSC+ approaches perform well qualitatively and even quantitatively within about $2\%$ In particular, they capture a small low-temperature drop of double occupancy in the renormalized classical regime.
@@
Figure 1: Double occupancy as function of the temperature $T$ for $n=1$ and $U=2$. Results for both TPSC+ and TPSC+Classical approaches are in good qualitative and quantitative agreement with DiagMC [Schaefer2021] results. Importantly, they capture a small rise and then drop of the double occupancy as $T$ enters the renormalized-classical regime
III.1.2 Correlation Length
Figure 2 shows the spin correlation length $\xi(\beta=1/T)$ defined by  [gauvin2023improved]
@@ III.1.2 Correlation Length
In addition to the TPSC+, TPSC+Classical approaches and DiagMC results, there are also results for the $D\Gamma A$ approximation [Schaefer2021]. The $D\Gamma A$ approximation has data for lower temperatures than DiagMC. Both TPSC+ approaches are in good agreement with the DiagMC and the $D\Gamma A$ approaches at high temperatures (small $\beta$). As temperature decrease, the correlation length enters an exponential regime. The onset of this regime $T_{exp},$ occurs at higher temperature for TPSC+ than for $D\Gamma A$. The lowest $T_{exp}$ is for TPSC+Classical but the difference with TPSC+ is small.
@@
Figure 2: Spin correlation length from TPSC+ and TPSC+Classical approaches compared to the DiagMC benchmark and $D\Gamma A$ approximation. The results are obtained as a function of the inverse temperature $\beta=1/T$ for the half-filled 2D Hubbard model with U = 2.
III.1.3 Self Energy
@@
Figure 3 shows the imaginary part of the self-energy as a function of Matsubara frequency at half-filling $n=1$, and interaction $U=2$ for three temperatures $T=0.0625,0.0654,0.1$ and nodal (N) $\mathbf{k}_{N}=(\pi/2,\pi/2)$ and antinodal (AN) $\mathbf{k}_{AN}=(\pi,0)$ points on the Fermi surface. The self-energy was calculated using the full expression for the second level approximation Eq. (8) with $\tilde{g}=1/4$ for TPSC+Classical and $\tilde{g}=3/8$ for TPSC+. The drop of the self-energy at the smallest Matsubara frequency is due to the pseudogap. This was used as a criterion for the pseudogap in Ref. [Schaefer2021].
@@
According to the DiagMC results [Schaefer2021] in Matsubara frequency, the crossover temperatures to the pseudogap regime (pseudogap) are $T=0.0654$ for the AN point and $T=0.0625$ for the N point. However, as we will see in the next section, the real-frequency results show a pseudogap at higher temperatures i.e. before $|\Sigma(k_{F},ik_{0})|$ becomes larger than $|\Sigma(k_{F},ik_{1})|$. The TPSC+Classical results with $\tilde{g}=1/4$ are in quantitative agreement with DiagMC results [Schaefer2021]. The TPSC+, which uses $\tilde{g}=3/8$, captures the pseudogap effect qualitatively but overestimates it quantitatively. In the rest of the paper we display results for TPSC+Classical with $\tilde{g}=1/4$.
@@
Figure 3: The imaginary part of the self-energy as a function of Matsubara frequencies at half-filling $n=1$, and interaction $U=2$ for three temperatures $T=0.0625,0.0654,0.1$ and nodal (N) $\mathbf{k}_{N}=(\pi/2,\pi/2)$ and antinodal (AN) $\mathbf{k}_{AN}=(\pi,0)$ points on the Fermi surface. The symbols are DiagMC results [Schaefer2021]. The drop of the self-energy at the smallest Matsubara frequency is due to the pseudogap.
@@
Figure 4 shows the absolute value of the self-energy at the lowest Matsubara frequency $-\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ as a function of temperature for the nodal and antinodal points on the Fermi surface. The symbols are data from DiagMC [Schaefer2021] and the lines represent results of the TPSC+Classical calculations. Agreement is very good. TPSC+Classical captures quantitatively the behavior in both the metallic ($T$ above minimum) and the pseudogap regimes ($T$ below minimum).
@@
Figure  4 presents the clearest evidence of the pseudogap in the Matsubara representation because this quantity is expected to diverge as $T\rightarrow 0$ (see Eq. (17) below). Indeed, the figure shows this tendency as temperature decreases. Note also that the absolute value of the self-energy $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ is larger for the antinodal point than for the nodal point. For the pseudogap regime this is expected from the asymptotic behavior of $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ for the antinodal and the nodal points. The asymptotic behavior of $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ is given by the classical contribution to the self-energy [Vilk1997], [vilk2023criteria]: For the antinodal point:
@@
III.1.4 Zero-Temperature Gap
The above expressions suggest a way to find the zero temperature gap $\Delta_{0}$. We plot the quantity $ik_{0}\Sigma(\mathbf{k},ik_{0})$ for the nodal and the antinodal points as a function of temperature. The temperature where both lines converge to the same asymptote is the temperature at which we can extract the value of $\Delta_{0}^{2}=ik_{0}\Sigma(\mathbf{k},ik_{0})$. Figure 5 illustrates the above approach. The lines for the nodal and antinodal points converge as temperature decreases when $T$ reaches $T\sim 0.005$. From this we obtained the value $\Delta_{0}\approx 0.15$.
@@ III.1.4 Zero-Temperature Gap
It is interesting to compare this result with the renormalized mean-field one, obtained from the generalized Stoner criterion Eq. (11). Solving Eq. (14), we obtain $\Delta_{0\_MF}\approx 0.18$. This is about $20\%$ larger than the result $\Delta_{0}\approx 0.15$ obtained using the above procedure ($\Delta_{0}^{2}=ik_{0}\Sigma(\mathbf{k},ik_{0})$ as $T\rightarrow 0$). Our analysis indicates that this difference is mostly because we used different values for factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=1/4$. Choosing the same value for both factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=3/8$ gives the gap value within $1\%$ of the mean field result $\Delta_{0\_MF}\approx 0.18$. We used the combination of factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=1/4$ because it gives the best fit of the Monte Carlo data for both spin susceptibility and the final self-energy. Since we have a good fit of the self-energy $\Sigma(\mathbf{k},ik_{0})$ to the Monte Carlo data 4, we believe that the value of $\Delta_{0}=0.15$ extracted from the above self-energy analysis is more accurate than the renormalized mean-field result. The mean-field gap Eq. (11) was derived using classical critical fluctuations only and the deviation from it gives a measure of quantum effects.
@@ III.2 Analytic continuation to real frequencies
At temperature $T=0.1$ (upper panel) a single quasi-particle peak in the spectral function $A(\mathbf{k},\omega)$ exists for regular points on the Fermi surface ($v_{F}\neq 0$), but for the Van Hove point $\mathbf{k}_{AN}=(\pi,0)$ the pseudogap is already opening up. Note that this is significantly higher than the crossover temperature $T=0.0654$ identified in Ref. [Schaefer2021] based on Matsubara frequency data only. The criterion for the pseudogap used in [Schaefer2021] was $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$. However, the real frequency data provides a different perspective and we conclude that the condition $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$ is a sufficient but not a necessary condition for the pseudogap. We also verified that for the nodal point the crossover to the pseudogap in $A(\mathbf{k}_{N},\omega)$ occurs at the temperature $T_{N}=0.08$. This is also higher than the crossover $T_{N}$ identified based on the above Matsubara-frequency criterion ($T_{N}=0.0625$). The difference between crossover temperatures for antinodal point $T_{AN}$ and nodal point $T_{N}$ is about $(T_{AN}-T_{N})\approx 0.02$, which is much larger than the difference based on the Matsubara criterion $(T_{AN}-T_{N})\approx 0.003$.
@@
Finally, at $T=0.005$ (bottom panel) the pseudogap is hard to distinguish, visually, from a real antiferromagnetic gap in the spectral function. The results for $T=0.005<\Delta$ give insight in the behavior at $T\rightarrow 0$. In addition, to quasiparticle peaks at $\omega=\pm\Delta$ and the pseudogap between them, there is also an incoherent background for $|\omega|>\Delta$. This is consistent with the results for the half-filled Hubbard model obtained using the non-linear sigma ($NL\sigma M$) description of the spin fluctuations [borejsza2003antiferromagnetism]. The latter work also shows that the peak at $|\omega|\rightarrow\Delta-0$ becomes infinitely narrow and infinitely high as $T\rightarrow 0$. On the other hand, the height of $A(\mathbf{k}_{F},\omega)$ remains finite when $|\omega|\rightarrow\Delta+0$. This behavior implies that the imaginary part of the self-energy $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ in [borejsza2003antiferromagnetism] is a step function around $|\omega|=\Delta$ with zero value inside the region $0<|\omega|<\Delta$ and a finite value outside the gap (at exactly $\omega=0$ there is negative delta function for $T\rightarrow 0$) . A different perspective on the behavior of the self-energy in the ordered state is provided in Ref. [sangiovanni2006static] in which the self-energy is continuous at $|\omega|=\Delta$ and goes to zero as $T\rightarrow 0$. The results in Ref. [sangiovanni2006static] are presented for the strong coupling limit and were obtained using dynamical mean-field theory (DMFT) in infinite dimension.
@@
Our result for $|\omega|<\Delta$ and $T\rightarrow 0$ is somewhat different in this respect from the results in [borejsza2003antiferromagnetism] and [sangiovanni2006static]. Specifically, the weight of $A(\mathbf{k}_{F},\omega)$ for $|\omega|<\Delta$ is very small but it is not zero. This is because in the one-loop like approximation that we use (Eq. (8), the quantum contribution to $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ is continuous and is a linear function at $|\omega|>>T$ (see bellow). Although it is small, it does not go to zero as $T\rightarrow 0$
@@
Figure 7 displays $\Sigma^{\prime\prime}(\mathbf{k}_{F}=\mathbf{k}_{AN},\omega)$ in our theory for $T=0.005$ at the antinode. There is a delta-function-like peak at $\omega=0$ (most of the peak is outside the picture) and a linear dependence of $\Sigma^{\prime\prime}(\mathbf{k}_{F}=\mathbf{k}_{AN},\omega)$ for $\omega>\pi T$. The former is due to the critical classical fluctuations and is responsible for the pseudogap. The latter is due to quantum fluctuations. It is linear in $\omega$ due to nesting, it is unrelated to the antiferromagnetic peak in the spin susceptibility and is qualitatively similar to the second order perturbation theory result [Lemay2000, Schaefer2021]. We checked explicitly that this quantum linear behavior of $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ for $\omega>\pi T$ is not significantly affected by temperature. Since $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ is finite at $|\omega|=\Delta\approx 0.15$ the peaks at $|\omega|\rightarrow\Delta-0$ remains finite as $T\rightarrow 0$. We think that this behavior is likely an artifact of the one-loop-like approximation that we used. The alternative is that the spectral function is not continuous at $T=0$ and that the weight inside the gap disappears suddenly at the $T=0$ phase transition.
@@
Figure 6: Evolution of the spectral function $A(\mathbf{k},\omega)$ with temperature. The results are for tree points on the Fermi surface. The parameters are $U=2$ and $n=1$. The upper panel shows that at the antinodal point $\mathbf{k}_{AN}=(\pi,0)$ ($v_{F}=0$) the pseudogap has started to develop for $T=0.1$. This is significantly higher than the pseudogap crossover temperature identified earlier [Schaefer2021] based on Matsubara data only. For regular points on the Fermi surface ($v_{F}\neq 0$) there is still a single particle peak in $A(\mathbf{k},\omega)$ at this temperature. The middle panel $T=0.0625$ shows the pseudogap in the spectral function at all points on the Fermi surface. The lower panel shows a well developed pseudogap at the very low temperature $T=0.005<\Delta$ that is hard, visually, to distinguish from a real antiferromagnetic gap.
@@
Figure 8 shows the evolution of the density of states $N(\omega)$ with temperature. We can see that at higher temperatures there a is a maximum at $\omega=0$. This maximum is due to the Van Hove singularity that in 2D leads to the divergence of $N(\omega)$ at $\omega=0$ in the non-interacting system. Note that at $T=0.0625$ the pseudogap is already developed in the spectral function $A(\mathbf{k}_{F},\omega)$ on the Fermi surface but it is not present in the density of states. This is because that quantity is an integral over all wave vectors, so until the pseudogap is well developed, its Fermi-surface contribution to the total density of states may be negligible. The pseudogap in the density of states eventually develops at lower temperatures and can be clearly seen in the lower panel ($T=0.005$) of Fig. 8.

### v2

@@ III.1.1 Spin Susceptibility
The TPSC+ study in the Ref. [56] has results for various fillings, temperatures and interaction strength for both two and one particle properties. The results are for TPSC+ and TPSC+SFM but, as we mentioned above, the results for $U\leq 4$ are quantitatively similar to results for TPSC+Classical with $\tilde{g}_{fb}=3/8$ and $\tilde{g}_{fb}=1/4$, respectively. For technical reasons, study in the Ref. [56] was limited to temperatures $T\geq 0.79$. In the current study, we were able to resolve convergence problem at low temperatures using methods that are explained in Appendix A. In the rest of the paper, we focus on low temperatures at half-filling, on the antiferromagnetic pseudogap regime and on estimates of the zero-temperature gap. We choose $U=2$ because this allows us to make quantitative comparison to the DiagMC results. We also found that at very low temperatures (large correlation length) the TPSC+Classical with $\tilde{g}_{fb}=3/8$ works the best and we use this value in the rest of the paper.
Figure 1: Maximal value of the spin susceptibility obtained from various TPSC+ approaches (see text) as a function of $U$ compared with Monte Carlo data [65]. Results are shown for $n=0.875$ and $T=0.2$. All TPSC+ approaches agree reasonably well with the benchmark Monte Carlo data for $U<=2.5$ at this temperature. For $U>4$ the Monte Carlo values of $\chi_{sp}(\mathbf{q}_{max},0)$ start to saturate and then decrease as a function of $U$. The latter behavior indicates transition to Mott physics. All TPSC+ approaches miss this behavior quantitatively but they do show a broad maximum for $U\geq 6$. This maximum is due to the effect of the feedback self-energy on the spin susceptibility.
III.1.2 Double Occupancy
@@
Figure 2 shows double occupancy $D$ as a function of temperature $T$. The results are for TPSC+, TPSC+Classical and for the DiagMC [52] benchmark results. The factor $\tilde{g}_{fb}$ for the feedback self-energy used in the calculation of the partially-dressed susceptibility $\chi^{(2)}$ was set to $\tilde{g}_{fb}=3/8$. Both TPSC+ approaches perform well qualitatively and even quantitatively within about $2\%$ In particular, they capture a small low-temperature drop of double occupancy in the renormalized classical regime.
@@
Figure 2: Double occupancy as a function of the temperature $T$ for $n=1$ and $U=2$. Results for both TPSC+ and TPSC+Classical approaches are in good qualitative and quantitative agreement with DiagMC [52] results. Importantly, they capture a small rise and then drop of the double occupancy as $T$ enters the renormalized-classical regime
III.1.3 Correlation Length
Figure 3 shows the spin correlation length $\xi(\beta=1/T)$ defined by  [56]
@@ III.1.3 Correlation Length
In addition to the TPSC+, TPSC+Classical approaches and DiagMC results, there are also results for the $D\Gamma A$ approximation [52]. The $D\Gamma A$ approximation has data for lower temperatures than DiagMC. Both TPSC+ approaches are in good agreement with the DiagMC and the $D\Gamma A$ approaches at high temperatures (small $\beta$). As temperature decrease, the correlation length enters an exponential regime. The onset of this regime $T_{exp},$ occurs at higher temperature for TPSC+ than for $D\Gamma A$. The lowest $T_{exp}$ is for TPSC+Classical but the difference with TPSC+ is small.
@@
Figure 3: Spin correlation length from TPSC+ and TPSC+Classical approaches compared with the DiagMC benchmark and $D\Gamma A$ approximation. The results are obtained as a function of the inverse temperature $\beta=1/T$ for the half-filled 2D Hubbard model with U = 2.
III.1.4 Self Energy
@@
Figure 4 shows the imaginary part of the self-energy as a function of Matsubara frequency at half-filling $n=1$, and interaction $U=2$ for three temperatures $T=0.0625,0.0654,0.1$ and nodal (N) $\mathbf{k}_{N}=(\pi/2,\pi/2)$ and antinodal (AN) $\mathbf{k}_{AN}=(\pi,0)$ points on the Fermi surface. The self-energy was calculated using the full expression for the second level approximation Eq. (8) with $\tilde{g}=1/4$ for TPSC+Classical and $\tilde{g}=3/8$ for TPSC+. The drop of the self-energy at the smallest Matsubara frequency is due to the pseudogap. This was used as a criterion for the pseudogap in Ref. [52].
@@
According to the DiagMC results [52] in Matsubara frequency, the crossover temperatures to the pseudogap regime (pseudogap) are $T=0.0654$ for the AN point and $T=0.0625$ for the N point. However, as we see in the next section, the real-frequency results show a pseudogap at higher temperatures i.e. before $|\Sigma(k_{F},ik_{0})|$ becomes larger than $|\Sigma(k_{F},ik_{1})|$. The TPSC+Classical results with $\tilde{g}=1/4$ are in quantitative agreement with DiagMC results [52]. The TPSC+, which uses $\tilde{g}=3/8$, captures the pseudogap effect qualitatively but overestimates it quantitatively. In the rest of the paper we display results for TPSC+Classical with $\tilde{g}=1/4$.
@@
Figure 4: The imaginary part of the self-energy as a function of Matsubara frequencies at half-filling $n=1$, and interaction $U=2$ for three temperatures $T=0.0625,0.0654,0.1$ and nodal (N) $\mathbf{k}_{N}=(\pi/2,\pi/2)$ and antinodal (AN) $\mathbf{k}_{AN}=(\pi,0)$ points on the Fermi surface. The symbols are DiagMC results [52]. The drop of the self-energy at the smallest Matsubara frequency is due to the pseudogap.
@@
Figure 5 shows the absolute value of the self-energy at the lowest Matsubara frequency $-\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ as a function of temperature for the nodal and antinodal points on the Fermi surface. The symbols are data from DiagMC [52] and the lines represent results of the TPSC+Classical calculations. Agreement is very good. TPSC+Classical captures quantitatively the behavior in both the metallic ($T$ above minimum) and the pseudogap regimes ($T$ below minimum).
@@
Figure  5 presents the clearest evidence of the pseudogap in the Matsubara representation because this quantity is expected to diverge as $T\rightarrow 0$ (see Eq. (17) below). Indeed, the figure shows this tendency as temperature decreases. Note also that the absolute value of the self-energy $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ is larger for the antinodal point than for the nodal point. For the pseudogap regime this is expected from the asymptotic behavior of $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ for the antinodal and the nodal points. The asymptotic behavior of $\mathrm{Im}\Sigma(\mathbf{k},ik_{0})$ is given by the classical contribution to the self-energy [10], [59]: For the antinodal point:
@@
III.1.5 Zero-Temperature Gap
The above expressions suggest a way to find the zero-temperature gap $\Delta_{0}$. We plot the quantity $ik_{0}\Sigma(\mathbf{k},ik_{0})$ for the nodal and the antinodal points as a function of temperature. The temperature where both lines converge to the same asymptote is the temperature at which we can extract the value of $\Delta_{0}^{2}=ik_{0}\Sigma(\mathbf{k},ik_{0})$. Figure 6 illustrates the above approach. The lines for the nodal and antinodal points converge as temperature decreases when $T$ reaches $T\sim 0.005$. From this we obtained the value $\Delta_{0}\approx 0.15$.
@@ III.1.5 Zero-Temperature Gap
It is interesting to compare this result with the renormalized mean-field one, obtained from the generalized Stoner criterion Eq. (11). Solving Eq. (14), we obtain $\Delta_{0\_MF}\approx 0.18$. This is about $20\%$ larger than the result $\Delta_{0}\approx 0.15$ obtained using the above procedure ($\Delta_{0}^{2}=ik_{0}\Sigma(\mathbf{k},ik_{0})$ as $T\rightarrow 0$). Our analysis indicates that this difference is mostly because we used different values for factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=1/4$. Choosing the same value for both factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=3/8$ gives the gap value within $1\%$ of the mean field result $\Delta_{0\_MF}\approx 0.18$. We used the combination of factors $\tilde{g}_{fb}=3/8$ and $\tilde{g}=1/4$ because it gives the best fit of the Monte Carlo data for both spin susceptibility and the final self-energy. Since we have a good fit of the self-energy $\Sigma(\mathbf{k},ik_{0})$ to the Monte Carlo data 5, we believe that the value of $\Delta_{0}=0.15$ extracted from the above self-energy analysis is more accurate than the renormalized mean-field result. The mean-field gap Eq. (11) was derived using classical critical fluctuations only and the deviation from it gives a measure of quantum effects.
@@ III.2 Analytic continuation to real frequencies
At temperature $T=0.1$ (upper panel) a single quasi-particle peak in the spectral function $A(\mathbf{k},\omega)$ exists for regular points on the Fermi surface ($v_{F}\neq 0$), but for the Van Hove point $\mathbf{k}_{AN}=(\pi,0)$ the pseudogap is already opening up. Note that this is significantly higher than the crossover temperature $T=0.0654$ identified in Ref. [52] based on Matsubara frequency data only. The criterion for the pseudogap used in [52] was $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$. However, the real frequency data provides a different perspective and we conclude that the condition $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$ is a sufficient but not a necessary condition for the pseudogap. In the case we study, this criterion gives a lower bound. We also verified that for the nodal point the crossover to the pseudogap in $A(\mathbf{k}_{N},\omega)$ occurs at the temperature $T_{N}=0.08$. This is also higher than the crossover $T_{N}$ identified based on the above Matsubara-frequency criterion ($T_{N}=0.0625$). The difference between crossover temperatures for antinodal point $T_{AN}$ and nodal point $T_{N}$ is about $(T_{AN}-T_{N})\approx 0.02$, which is much larger than the difference based on the Matsubara criterion $(T_{AN}-T_{N})\approx 0.003$.
@@
Finally, at $T=0.005$ (bottom panel) the pseudogap is hard to distinguish, visually, from a real antiferromagnetic gap in the spectral function. The results for $T=0.005<\Delta$ give insight in the behavior at $T\rightarrow 0$. In addition, to quasiparticle peaks at $\omega=\pm\Delta$ and the pseudogap between them, there is also an incoherent background for $|\omega|>\Delta$. This is consistent with the results for the half-filled Hubbard model obtained using the nonlinear sigma ($NL\sigma M$) description of the spin fluctuations [66]. The latter work also shows that the peak at $|\omega|\rightarrow\Delta-0$ becomes infinitely narrow and infinitely high as $T\rightarrow 0$. On the other hand, the height of $A(\mathbf{k}_{F},\omega)$ remains finite when $|\omega|\rightarrow\Delta+0$. This behavior implies that the imaginary part of the self-energy $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ in [66] is a step function around $|\omega|=\Delta$ with zero value inside the region $0<|\omega|<\Delta$ and a finite value outside the gap (at exactly $\omega=0$ there is negative delta function for $T\rightarrow 0$) . A different perspective on the behavior of the self-energy in the ordered state is provided in Ref. [67] in which the self-energy is continuous at $|\omega|=\Delta$ and goes to zero as $T\rightarrow 0$. The results in Ref. [67] are presented for the strong coupling limit and were obtained using dynamical mean-field theory (DMFT) in infinite dimension.
@@
Our result for $|\omega|<\Delta$ and $T\rightarrow 0$ is somewhat different in this respect from the results in [66] and [67]. Specifically, the weight of $A(\mathbf{k}_{F},\omega)$ for $|\omega|<\Delta$ is very small but it is not zero. This is because in the one-loop like approximation that we use (Eq. (8), the quantum contribution to $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ is continuous and is a linear function at $|\omega|>>T$ (see bellow). Although it is small, it does not go to zero as $T\rightarrow 0$
@@
Figure 8 displays $\Sigma^{\prime\prime}(\mathbf{k}_{F}=\mathbf{k}_{AN},\omega)$ in our theory for $T=0.005$ at the antinode. There is a delta-function-like peak at $\omega=0$ (most of the peak is outside the picture) and a linear dependence of $\Sigma^{\prime\prime}(\mathbf{k}_{F}=\mathbf{k}_{AN},\omega)$ for $\omega>\pi T$. The former is due to the critical classical fluctuations and is responsible for the pseudogap. The latter is due to quantum fluctuations. It is linear in $\omega$ due to nesting, it is unrelated to the antiferromagnetic peak in the spin susceptibility and is qualitatively similar to the second order perturbation theory result [68, 52]. We checked explicitly that this quantum linear behavior of $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ for $\omega>\pi T$ is not significantly affected by temperature. Since $\Sigma^{\prime\prime}(\mathbf{k}_{F},\omega)$ is finite at $|\omega|=\Delta\approx 0.15$ the peaks at $|\omega|\rightarrow\Delta-0$ remains finite as $T\rightarrow 0$. We think that this behavior is likely an artifact of the one-loop-like approximation that we used. The alternative is that the spectral function is not continuous at $T=0$ and that the weight inside the gap disappears suddenly at the $T=0$ phase transition.
@@
Figure 7: Evolution of the spectral function $A(\mathbf{k},\omega)$ with temperature. The results are for tree points on the Fermi surface. The parameters are $U=2$ and $n=1$. The upper panel shows that at the antinodal point $\mathbf{k}_{AN}=(\pi,0)$ ($v_{F}=0$) the pseudogap has started to develop for $T=0.1$. This is significantly higher than the pseudogap crossover temperature identified earlier [52] based on Matsubara data only. For regular points on the Fermi surface ($v_{F}\neq 0$) there is still a single particle peak in $A(\mathbf{k},\omega)$ at this temperature. The middle panel $T=0.0625$ shows the pseudogap in the spectral function at all points on the Fermi surface. The lower panel shows a well developed pseudogap at the very low temperature $T=0.005<\Delta$ that is hard, visually, to distinguish from a real antiferromagnetic gap.
@@
Figure 9 shows the evolution of the density of states $N(\omega)$ with temperature. We can see that at higher temperatures there a is a maximum at $\omega=0$. This maximum is due to the Van Hove singularity that in 2D leads to the divergence of $N(\omega)$ at $\omega=0$ in the noninteracting system. Note that at $T=0.0625$ the pseudogap is already developed in the spectral function $A(\mathbf{k}_{F},\omega)$ on the Fermi surface but it is not present in the density of states. This is because that quantity is an integral over all wave vectors, so until the pseudogap is well developed, its Fermi-surface contribution to the total density of states may be negligible. The pseudogap in the density of states eventually develops at lower temperatures and can be clearly seen in the lower panel ($T=0.005$) of Fig. 9.